[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Sun Jul 16, 2017

Christopher Mello - NOAA Federal christopher.mello at noaa.gov
Sun Jul 16 14:01:29 UTC 2017


All,
NAWIPS should return to normal. I removed old files from the satellite
partition which was full. This caused the data dropouts. We are now  at 74
percent which is very low. I wrote a scour script that removed all of the
GOES-R data that was from the July 14 and July 15. This should keep us
going until Monday when we  will fix the baseline  scouring programs.


Chris


On Sun, Jul 16, 2017 at 5:23 AM, HSU Shift Log <nhc.tsbadmin at noaa.gov>
wrote:

> *HSU Shift Log for Sun Jul 16 2017 05:22:57 GMT-0400 (EDT)*
>
> *Shift:* A
>
> *Forecaster:* S R Stewart
>
> *Technical Issues Summary:*
>
> *ATCF issues:* None.
>
> *N-AWIPS issues:* None.
>
> *AWIPS-2/NCP issues:* None.
>
> *Other Techincal issues:* Lots of GOES-16 IR data dropouts on shift;
> however, no problems were noted with the other channels.
>
> *Miscellaneous issues:* None.




-- 
Chris Mello
National Hurricane Center
TSB
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170716/904bf09d/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list